### v1

# Resonant Photon-Axion Mixing Driven by Dark Matter Oscillations

Abstract
Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem. We show that photon propagation through an axion dark matter background in the presence of a magnetic field is governed by Floquet physics, distinct from conventional static or adiabatic mixing paradigms. Coherent photon–axion mode conversion occurs when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$) with the axion mass $m_{a}$, even far from the standard level-crossing condition $\Delta_{\gamma}\approx\Delta_{a}$. Crucially, this resonance disappears entirely if the axion oscillations are averaged over, and is therefore systematically missed in conventional static or adiabatic treatments. This driven resonance represents a unitary Floquet mode-mixing process and is fundamentally distinct from parametric instability or stimulated axion decay, preserving the axion dark matter number density. We develop a general Floquet framework for photon propagation in oscillating axion backgrounds, revealing that resonant mixing generates robust polarization effects during propagation. As an astrophysical application, we apply this mechanism to the realistic environment of the blazar 3C 279 to derive concrete constraints on the axion–photon coupling. While the observational manifestation depends on environmental conditions, the underlying driven mixing mechanism is generic to coherent axion dark matter, revealing a previously overlooked regime of photon–axion conversion.
Introduction.– The propagation of waves in time-dependent media is a central problem across physics. When the properties of a medium vary periodically in time, the resulting dynamics are generically governed by Floquet physics, leading to resonant mode mixing that cannot be captured by static or adiabatic approximations. Such driven propagation phenomena are well known in condensed matter and quantum optics [1965PhRv..138..979S, Sambe:1973cnm, 2004PhR...390....1C, Oka:2009djg, Eckardt:2016lof], yet they have received little attention in the context of particle astrophysics.
A prominent and well-motivated realization of a coherently oscillating medium arises if dark matter (DM) [Zwicky:1933gu, Bertone:2018krk] is composed of axions [Peccei:1977hh, Peccei:1977ur, Weinberg:1977ma, Wilczek:1977pj, Cheng:1987gp] or axion-like particles [Preskill:1982cy, Abbott:1982af, Dine:1982ah, Arvanitaki:2009fg, Graham:2015ouw, Peccei:2006as, Marsh:2015xka, DiLuzio:2020wdo, Adams:2022pbo, OHare:2024nmr]. In this case, the DM forms a classical field that oscillates at a frequency set by the axion mass over macroscopic coherence scales. Photons propagating through this background therefore experience a medium whose optical properties are intrinsically periodic in time. Nevertheless, conventional treatments of photon–axion mixing [Raffelt:1987im, Raffelt:1996wa] typically average over the axion oscillations or assume an effectively static background, thereby reducing the problem to a time-independent level-crossing or adiabatic conversion scenario.
In astrophysical environments, the photon–axion system is typically analyzed in two regimes: photon–axion conversion (the Primakoff effect) [Primakoff:1951iae, Sikivie:1983ip, Raffelt:1987im] and rotation of the polarization plane (birefringence) [Carroll:1989vb, Carroll:1998zi, Harari:1992ea]. Conventional analyses employ adiabatic or perturbative approximations, effectively averaging over rapid axion oscillations or treating the background as quasi-static.
In this work, we show that these assumptions miss a qualitatively novel class of resonant propagation effects. We demonstrate that photon–axion mixing in a coherent axion DM background is fundamentally a periodically driven problem, governed by Floquet dynamics rather than static dispersion matching. Our analysis reveals that the intrinsic time periodicity of the axion background acts as a parametric driver, actively inducing coherent mixing between photon and axion modes. This process leads to parametric resonance when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion mass as $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$). Consequently, resonant conversion can occur far from the conventional level-crossing resonance condition $\Delta_{\gamma}\approx\Delta_{a}$ [Pshirkov:2007st, Hook:2018iia]. This resonance arises purely from temporal phase matching induced by the oscillating background and does not rely on spatial inhomogeneities or fine-tuned environmental profiles. The driven resonance is not a higher-order correction to static mixing: it relies on the coherent axion oscillation and vanishes completely once this time dependence is averaged out.
The driven resonance identified here is fundamentally distinct from parametric instability or stimulated axion decay ($m_{a}\simeq 2\omega$) [Tkachev:2014dpa, Arza:2019nta, Caputo:2018vmy, Caputo:2018ljp]. The propagation dynamics remain unitary, the axion background number density is conserved, and the system exhibits coherent Floquet mode mixing rather than exponential amplification.
We develop a general Floquet framework for photon propagation in oscillating axion backgrounds and show that this mechanism leads to robust and generic consequences for photon polarization during propagation. While specific observational manifestations depend on environmental conditions, the underlying driven mixing phenomenon is intrinsic to coherent axion DM and represents a previously overlooked regime of photon–axion conversion. Our results establish a systematic framework for analyzing propagation effects in time-dependent backgrounds, revealing previously unexplored resonant regimes.
Driven photon–axion Mixing.– Consider photon propagation in the presence of a magnetic field and a coherently oscillating axion background (see also [Espriu:2014lma, Masaki:2019ggg, McDonald:2019wou] for a similar system),
$\bar{a}(t,z)=m_{a}^{-1}\sqrt{2\rho_{\text{DM}}}\cos(m_{a}t+\varphi_{a})$ (1)
where $m_{a}$ is the axion mass, $\rho_{\text{DM}}$ is the DM density, and $\varphi_{a}$ is the phase of the axion background. In an appropriate polarization basis, the propagation equation takes the form
$i\partial_{z}\Psi=\left[\mathcal{H}_{0}+\mathcal{H}_{1}\right]\Psi$ (2)
where the state vector $\Psi$ includes the two photon polarization states and the axion field. The photon–axion system is governed by a Hermitian mixing Hamiltonian whose coefficients contain an explicit periodic modulation at frequency $m_{a}$, induced by the axion background. $\mathcal{H}_{0}$ incorporates the usual photon dispersion $\Delta_{\perp}$ and $\Delta_{\parallel}$, axion mass term $\Delta_{a}$, and magnetic-field-induced mixing $\Delta_{\mathrm{B}}$, and is given by
$\mathcal{H}_{0}=-\begin{pmatrix}\Delta_{\perp}&0&0\\ 0&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$. (3)
$\mathcal{H}_{1}$ encodes the axion-induced birefringence
$\mathcal{H}_{1}(z)=-\begin{pmatrix}0&\Delta_{F}&0\\ \Delta_{F}^{*}&0&0\\ 0&0&0\end{pmatrix}$, (4)
with
$\Delta_{F}=ig_{a\gamma}(\partial_{t}{\bar{a}}+\partial_{z}\bar{a})/2.$ (5)
In the linear polarization basis, the axion-induced Faraday term is purely imaginary, reflecting its interpretation as a rotation generator. The Hamiltonian is manifestly Hermitian, and its only explicit time dependence arises from the oscillating axion background:
$\Delta_{F}=-i\Delta_{\text{DM}}\sin(m_{a}t+\varphi_{a}),$ (6)
where the amplitude is $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{\rho_{\text{DM}}/2}$. Along the photon trajectory we take the eikonal propagation approximation $t\simeq z$ (with $c=1$).
The propagation equation represents a three-level driven system with period $T=2\pi/m_{a}$. It is convenient to remove the explicit time dependence induced by the axion background by transforming to a co-rotating frame that follows the axion-induced phase modulation. This is achieved by a time-dependent unitary transformation that eliminates the oscillatory off-diagonal term at the expense of introducing harmonic couplings between propagation eigenstates.
We perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ to reallocate the rapid oscillations. Applying the rotation operator $U=u\oplus 1$, where $u$ acts on the photon subspace as
$u(z)=\exp\left[-i\theta(z)\sigma_{y}\right],$ (7)
the fast oscillating sinusoidal term is absorbed into a phase-dependent rotation angle $\theta(z)=\alpha\cos\phi(z)$ with $\alpha=\Delta_{\text{DM}}/m_{a}$ and $\phi(z)=m_{a}z+\varphi_{a}$. This rotation is chosen such that the term generated by $U^{\dagger}i\partial_{z}U$ exactly cancels the rapidly oscillating birefringence $\Delta_{F}$. Assuming negligible QED birefringence with $|\Delta_{\perp}-\Delta_{\parallel}|/2\ll m_{a}$ and $\Delta_{\mathrm{DM}}$, the effective Hamiltonian becomes:
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$, (8)
where $\Delta_{\gamma}=(\Delta_{\parallel}+\Delta_{\perp})/2$ is the mean photon refractive index. The unitary rotation is exact algebraically and does not change the quasienergy spectrum; it simply relocates the periodic dependence.
In the regime $\Delta_{\mathrm{B}}\ll|\Delta_{\gamma}-\Delta_{a}|$, the off-resonant photon mode can be adiabatically eliminated, allowing the dynamics to be analyzed by treating each photon mode-axion pair independently within the rotating-wave approximation. The resulting two-level systems are governed by the couplings $G^{(1)}=\Delta_{\mathrm{B}}\sin\theta$ and $G^{(2)}=\Delta_{\mathrm{B}}\cos\theta$. These couplings depend on $\theta(z)$; $\sin\theta$ and $\cos\theta$ can be written as linear combinations of $e^{\pm i\theta}$.
Using the Jacobi-Anger expansion, $e^{i\alpha\cos\phi}=\sum_{\ell}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}$ with the Bessel function $J_{\ell}(\alpha)$, we expand $G^{(1,2)}$ into Fourier harmonics. The couplings can be written as
$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z},$ (9)
where $g_{\ell}^{(1)}$ is nonzero only for odd $\ell$, while $g_{\ell}^{(2)}$ is nonzero only for even $\ell$. Explicitly, the $\ell$-th harmonic coupling scales as $g_{\ell}\sim\Delta_{\mathrm{B}}J_{\ell}(\alpha)$ up to phase factors, where the Bessel function $J_{\ell}(\alpha)$ behaves roughly as an oscillatory sine or cosine that decays proportionally to $\alpha^{-1/2}$.
In this co-rotating basis, the driven photon–axion system acquires a Floquet structure: the axion oscillation generates sidebands separated by integer multiples of $m_{a}$, and resonant mode mixing occurs when the propagation eigenvalue splitting matches one of these harmonics. In this basis, the axion-induced term acts as a periodic off-diagonal drive that couples propagation eigenstates separated in quasi-energy. Transitions between eigenstates become resonant whenever the energy mismatch $\Delta=\Delta_{\gamma}-\Delta_{a}$ is compensated by the absorption or emission of integer multiples of the driving frequency.
We remove the free propagation phases by defining $\Phi=\operatorname{diag}\left(e^{i\Delta_{\gamma}z},e^{i\Delta_{a}z}\right)\tilde{\Phi}_{\gamma}$. Under this transformation, the $\ell$-th Fourier component of the coupling acquires a phase factor $e^{i(\Delta_{\gamma}-\Delta_{a}-\ell m_{a})z}$, with the corresponding detuning given by $\Delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}$. Resonance occurs when the phase is stationary, yielding the Floquet resonance condition:
$\Delta_{\gamma}-\Delta_{a}\approx nm_{a},\quad n\in\mathbb{Z}.$ (10)
This condition characterizes coherent mode mixing within a time-periodic background, demonstrating that substantial conversion can occur even far from the standard level-crossing resonance $\Delta_{\gamma}=\Delta_{a}$.
Near a given resonance $n$, keeping only the corresponding Fourier mode amounts to a rotating-wave approximation. The two-state system undergoes Rabi oscillations with an effective frequency $\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\Delta_{n}/2)^{2}}$. The conversion probability is given by
$P_{\gamma\to a}\approx\frac{|g_{n}|^{2}}{\Omega_{\rm R}^{2}}\sin^{2}\big(\Omega_{\rm R}\,z\big).$ (11)
The resonance is maintained as long as the detuning satisfies $|\Delta_{n}(z)|\lesssim 2|g_{n}|$, a regime where the peak conversion efficiency remains at least 50%. Note that the photon state $\gamma$ here refers to the effective photon mode under rotation. It can be verified that in the static limit $m_{a}\to 0$, all sidebands collapse and the system reduces exactly to conventional photon–axion mixing. Additionally, for small modulation amplitude $\Delta_{\mathrm{DM}}\ll m_{a}$, only the lowest harmonics contribute, and the effect reduces to a single-frequency perturbation of the standard oscillation probability.
More detailed derivations are provided in the supplementary material. An equivalent description in the circular polarization basis is also discussed in the supplementary material, where the axion background appears as a periodic modulation of the refractive indices. Related effects of axion-induced frequency modulation have been discussed previously, primarily in the context of birefringence or modifications of the dispersion relation [McDonald:2019wou]. In those treatments, the time-dependent phase is typically treated perturbatively or averaged over, and the associated Floquet sidebands are not explicitly identified. As a result, the resonant mode mixing and the corresponding Rabi oscillations discussed here were not recognized.
We emphasize that the Floquet resonance described here is fundamentally different from parametric instability or stimulated axion decay with $m_{a}\simeq 2\omega$. Here, the DM number density remains conserved, with the oscillating background providing the necessary energy ”kick” to bridge the momentum mismatch between the photon and axion dispersion relations. The Floquet coherent mixing in a periodic background is also dynamically distinct from the phenomenon of parametric instability. While both involve frequency matching, the former constitutes a unitary driven two-level problem characterized by stable mode-switching. In contrast, parametric instability requires a non-Hermitian structure for energy pumping (typically associated with sum-frequency channels  [^1] In the context of stimulated axion decay, the instability of the photon–axion system—considering both axion DM and magnetic fields but omitting the plasma background—is explored in Ref. [Masaki:2019ggg], where an instability condition $\sqrt{k_{\gamma}^{2}+m_{a}^{2}}+k_{\gamma}=nm_{a}$ is empirically identified.), which is kinematically inaccessible in the difference-frequency channel. Consequently, our resonance analysis leads to coherent oscillations that preserve the background number density, rather than the exponential growth characteristic of unstable systems. More discussion of this distinction is provided in the supplementary material.
Representative Propagation.– To illustrate the physical implications of the Floquet resonance, we consider a representative photon propagation scenario in which the photon dispersion $\Delta_{\gamma}$ varies slowly along the propagation direction, while the axion background remains coherent over the relevant length scale. This setup captures the essential physics of resonance crossing without requiring detailed modeling of a specific astrophysical environment.
As the photons propagate, the slowly varying dispersion causes the mismatch $\Delta(z)=\Delta_{\gamma}(z)-\Delta_{a}$ to sweep across the Floquet resonance condition $\Delta=nm_{a}$. When this occurs, the driven mode coupling becomes resonant, leading to efficient photon–axion conversion. Importantly, this resonance depends on temporal phase matching rather than spatial level crossing, and therefore persists in the absence of sharp gradients.
Figure 1: Evolution of the photon–axion conversion probability $P_{\gamma\to a}$ along the jet of blazar. The blue solid line represents the case with an axion DM background, while the orange line shows the standard adiabatic conversion without DM. The vertical red dashed lines indicate the locations where the resonance condition $\Delta_{\gamma}(r)-\Delta_{a}\approx nm_{a}$ is satisfied.
Fig. 1 shows the photon–axion conversion probability $P_{\gamma\to a}$ as a function of propagation distance for a representative choice of parameters satisfying the Floquet resonance condition. A pronounced enhancement occurs when $\Delta(z)$ crosses $\Delta=m_{a}$, demonstrating coherent mode mixing induced by the axion-driven modulation. Away from resonance, the conversion probability remains small and oscillatory, consistent with conventional photon–axion mixing.
The resonant region acts as a phase-matched mixing layer, in which the axion oscillation continuously supplies the energy mismatch required for conversion. The resulting behavior is analogous to Rabi oscillations in driven two-level systems, with the conversion probability bounded by unitarity.
We have verified that the resonance persists under moderate variations of the dispersion profile and in the presence of stochastic fluctuations, provided that the axion field remains coherent over the resonance region. Thus, the effect is not a fine-tuned artifact of a specific background profile. Detailed numerical analysis is presented in the Supplementary Material.
A distinctive consequence of this driven mixing is the generation of polarization asymmetries between propagation modes. In particular, the stochastic crossing of Floquet resonances induces circular polarization with a characteristic variance set by the resonance strength. Since the axion-induced modulation affects the two circular polarization states with opposite signs, Floquet resonances generically break the symmetry between them. As a result, photon propagation through an axion DM background becomes intrinsically polarization-sensitive, even for initially unpolarized states.
In particular, repeated or stochastic crossings of Floquet resonances lead to a net circular polarization with vanishing mean but nonzero variance. This stochastic circular polarization arises from coherent, unitary mode mixing and is therefore qualitatively distinct from polarization generated by absorption or scattering processes. The magnitude of the variance is directly controlled by the resonance strength, thereby providing a direct probe of the driven mixing mechanism. Importantly, this polarization signature does not rely on fine-tuned source properties or detailed environmental modeling. It follows directly from the existence of Floquet resonances in photon propagation through a coherent axion background.
Constraints from the balazar 3C 279.– We validate this theoretical framework with full numerical solutions and apply it to the optical polarimetry of the blazar 3C 279 as an example. The parametric resonance breaks the symmetry between circular polarization states, inducing a stochastic signal in the blazar jet. Detailed configurations are provided in the supplementary material. To derive constraints on the axion–photon coupling $g_{a\gamma}$, we adopt a conservative statistical criterion: the standard deviation of the axion-induced circular polarization $\sigma_{\mathrm{std}}$ must not exceed the characteristic observational uncertainty of the blazar 3C 279 ($\sigma_{\Pi_{C}}\approx 0.30\%$ [Liodakis:2021els]). This ensures consistency between the predicted stochastic signal and the measured noise level. For a given astrophysical source, the axion phase $\varphi_{a}$ is effectively constant over observational timescales. However, since its value is a priori unknown, the axion-induced circular polarization behaves observationally as a stochastic contribution with zero mean. Even when the axion-induced circular polarization exhibits a bimodal distribution for certain masses, its sign and magnitude remain unpredictable due to the unknown axion phase, and the variance provides a conservative measure of its observational impact. Requiring its variance not to exceed the instrumental sensitivity therefore provides a conservative consistency bound. Fig. 2 shows the illustrative constraint on parameter space, with CAST bounds [CAST:2017uph] included for comparison.
For $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, we obtain conservative limits as $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$. The distinctive band-like structure of the exclusion contour is a signature of the Floquet parametric resonance mechanism, reflecting the discrete nature of the resonance condition. Two features emerge: (1) The conversion efficiency, affected by the Bessel functions $J_{n}(\Delta_{\text{DM}}/m_{a})$, is maximized when the effective momentum mismatch is bridged by integer harmonics of $m_{a}$. The sharp dips in the exclusion limit correspond to ”sweet spots” where this resonance condition is optimally satisfied within the jet profile. (2) As $m_{a}$ varies, the dominant contribution shifts between different resonance orders $n$. Weaker constraints between bands reflect off-resonance regimes, where the system approaches the adiabatic limit, resulting in reduced circular polarization production.
The constraints in Fig. 2 show that axion-induced parametric resonance leads to verifiable physical consequences rather than negligible perturbative effects. Although the current limit region is set by the specific noise floor of 3C 279, the concrete limit validates the physical relevance of our framework.
Figure 2: Exclusion contours on the axion–photon coupling $g_{a\gamma}$ derived from 3C 279 optical polarimetry. The shaded region indicates parameters excluded by the non-observation of excess circular polarization. The resulting exclusion exhibits a distinctive band structure that directly reflects the discrete Floquet resonance condition $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, providing a characteristic signature of the driven mixing mechanism.
Conclusions.– Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem. We have shown that photon propagation in axion DM therefore exhibits Floquet dynamics, leading to resonant photon–axion mode conversion beyond the conventional static or adiabatic mixing paradigm. This driven resonance arises when the dispersion mismatch between photons and axions is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, enabling coherent conversion even far from the standard level-crossing condition.
The resulting Floquet resonance represents a distinct class of propagation phenomena. Unlike traditional resonant conversion, which is localized to specific spatial regions where plasma and axion masses coincide, the driven resonance identified here originates from temporal phase matching induced by the oscillating DM background. As a consequence, resonant mixing can persist over extended propagation distances, allowing small conversion amplitudes to accumulate coherently and imprint robust signatures on photon polarization.
The significance of this mechanism depends on the stability of the driven system against environmental fluctuations. In rapidly varying media, strong gradients can disrupt the phase coherence required for Floquet mixing, suppressing the resonance. In sufficiently smooth environments, however, the driven propagation effect can dominate photon–axion conversion, opening a coherence-based window for probing axion dark matter that is complementary to localized, high-density resonance searches.
More broadly, the Floquet framework developed here is not restricted to a single axion species. Because the resonance condition relies on the periodic structure of the background rather than its microscopic identity, analogous driven conversion phenomena may arise in multi-axion systems or more complex dark-sector environments. Our results therefore establish a general foundation for studying resonant wave propagation in coherently oscillating media, revealing a previously unexplored regime of dark-matter-induced photon conversion.
Acknowledgements.
Acknowledgments.– The authors would like to thank the computing cluster of School of Fundamental Physics and Mathematical Sciences at Hangzhou Institute for Advanced Study, UCAS for providing the computational resources used in this work. This work is supported by the National Key Research and Development Program of China Grant No. 2020YFC2201502, and the National Natural Science Foundation of China under Grants No. 12475065, 12250010, 12447131, 12447105, and 12575113.
Supplementary Material for Resonant photon–axion Mixing
Driven by Dark Matter Oscillations
Run-Min Yao, Xiao-Jun Bi, Peng-Fei Yin, and Qing-Guo Huang
This Supplementary Material (SM) contains additional calculation and derivation in support of the results presented in this work. First, we give a full derivation of the mixing matrix $\mathcal{M}$. Then, we present additional details regarding the derivation of the two-level systems. Next, we discuss the relation between our work and parametric instability. Finally, we present a numerical analysis and apply the formalism to the relativistic jet of blazar 3C 279.
I photon–axion Conversion in an Axion Background
The Lagrangian for the photon field $A_{\mu}$ and axion field $a$ can be written as
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-A_{\mu}j^{\mu}+\frac{1}{2}(\partial_{\mu}a)^{2}-\frac{1}{2}m_{a}^{2}a^{2}-\frac{1}{4}g_{a\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu},$ (S1)
where $F_{\mu\nu}$ and $\tilde{F}^{\mu\nu}$ denote the electromagnetic field tensor and its dual tensor, respectively, $j^{\mu}=(\rho,\mathbf{J})$ is the electromagnetic current density, and $g_{a\gamma}$ is the axion–photon coupling. Adopting the temporal gauge $A_{0}=A^{0}=0$ and fixing the residual gauge freedom by imposing the Coulomb (radiation) gauge $\bm{\nabla}\cdot\mathbf{A}=0$, the equations of motion are:
$\partial_{t}^{2}\mathbf{A}-\nabla^{2}\mathbf{A}=\mathbf{J}+g_{a\gamma}\partial_{t}a\bm{\nabla}\times\mathbf{A}-g_{a\gamma}\bm{\nabla}a\times\partial_{t}\mathbf{A},$ (S2)
$-\bm{\nabla}\cdot\partial_{t}\mathbf{A}=\rho-g_{a\gamma}\bm{\nabla}a\cdot\left(\bm{\nabla}\times\mathbf{A}\right),$ (S3)
$\partial_{t}^{2}a-\nabla^{2}a+m_{a}^{2}a=-g_{a\gamma}\partial_{t}\mathbf{A}\cdot\left(\bm{\nabla}\times\mathbf{A}\right).$ (S4)
For a weakly magnetized cold plasma, the dielectric tensor is isotropic, $\bm{\epsilon}=(1-\frac{\omega_{\mathrm{pl}}^{2}}{\omega^{2}})\mathbb{I}$, where $\omega_{\mathrm{pl}}=\sqrt{4\pi\alpha n_{e}/m_{e}}$ is the plasma frequency and $\omega$ is the energy of the photon–axion system. In the radiation gauge, longitudinal plasma oscillations decouple from the propagating electromagnetic modes. In Fourier space, the induced plasma current therefore contributes an isotropic dispersion term $\mathbf{J}_{\mathrm{ind}}(\omega)=-\omega^{2}(\mathbb{I}-\bm{\epsilon})\mathbf{A}(\omega)=-\omega_{\mathrm{pl}}^{2}\mathbf{A}(\omega)$ to both photon polarizations, which can be absorbed into an effective photon mass.
We apply the linearized perturbation to the equation and write the fields into background and perturbation
$\mathbf{A}=\bar{\mathbf{A}}+\delta\mathbf{A},\qquad a=\bar{a}+\delta a,$ (S5)
where $\bar{a}$ and $\bar{\mathbf{B}}=\bm{\nabla}\times\bar{\mathbf{A}}$ are the background axion and magnetic fields. The perturbations satisfy the equations of motion:
$\partial_{t}^{2}\delta\mathbf{A}-\nabla^{2}\delta\mathbf{A}+\omega_{\mathrm{pl}}^{2}\delta\mathbf{A}=g_{a\gamma}\bar{\mathbf{B}}\partial_{t}\delta a+g_{a\gamma}\partial_{t}\bar{a}\bm{\nabla}\times\delta\mathbf{A}-g_{a\gamma}\bm{\nabla}\bar{a}\times\partial_{t}\delta\mathbf{A},$ (S6)
$-\bm{\nabla}\cdot\partial_{t}\delta\mathbf{A}=\delta\rho-g_{a\gamma}\bm{\nabla}\bar{a}\cdot\left(\bm{\nabla}\times\delta\mathbf{A}\right)-g_{a\gamma}\bm{\nabla}\delta a\cdot\bar{\mathbf{B}},$ (S7)
$\partial_{t}^{2}\delta a-\nabla^{2}\delta a+m_{a}^{2}\delta a=-g_{a\gamma}\bar{\mathbf{B}}\cdot\partial_{t}\delta\mathbf{A}.$ (S8)
For a magnetic field perpendicular to the photon propagation direction and an axion gradient parallel to it, the source terms in Eq.(S7) vanish. The last two terms in Eq. (S6) characterize the birefringence effect induced by the axion DM. The first terms on the RHS of Eq. (S6) and Eq. (S8) reflect the axion–photon couplings. In conventional analysis regarding the photon–axion conversion, the variation of the axion DM is neglected.
We choose the direction of the wave vector $\hat{\mathbf{k}}$ along the $z$-axis and adopt the plane-wave Ansatz
$\delta\mathbf{A}=-i\tilde{\mathbf{A}}(z)e^{-i\omega t},\qquad\delta a=\tilde{a}(z)e^{-i\omega t},$ (S9)
where $\delta\mathbf{A}$ includes an additional factor of i-i compared to $\delta a$. This choice ensure that the axion–photon coupling terms in the mixing matrix remain real, consistent with the standard form of the equations of motion found in most literature. Subsequently, we derive
[…]
$(\partial_{z}^{2}+\omega^{2}-m_{a}^{2})\tilde{a}=-g_{a\gamma}\bar{\mathbf{B}}\omega\tilde{\mathbf{A}}$. (S11)
The equations can be rearranged as
$(\partial_{z}^{2}+\omega^{2}+2\omega\mathcal{M})\psi=0,$ (S12)
[…]
$\mathcal{M}(z)=\begin{pmatrix}\Delta_{\perp}&\Delta_{F}&0\\ \Delta_{F}^{*}&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$, (S13)
where $\Delta_{\perp}=\Delta_{\parallel}\equiv-\omega_{\mathrm{pl}}^{2}/(2\omega)$, $\Delta_{a}\equiv-m_{a}^{2}/(2\omega)$, and $\Delta_{\mathrm{B}}\equiv g_{a\gamma}B/2$. The $\Delta_{F}$ term originates from the axion-gradient-induced birefringence and is formally analogous to the Faraday rotation term. Explicitly, it is given by
$\Delta_{F}\equiv\frac{g_{a\gamma}}{2\omega}(\partial_{t}\bar{a}\partial_{z}+i\omega\partial_{z}\bar{a})$ (S14)
where the operator $\partial_{z}$ in the first term acts on the photon field. Within the WKB approximation, $i\partial_{z}\psi\approx\omega\psi$, the operator $\partial_{z}$ acting on $\psi$ can be replaced by $i\omega$, yielding
$\Delta_{F}=\frac{i}{2}g_{a\gamma}(\partial_{t}\bar{a}+\partial_{z}\bar{a}).$ (S15)
The second-order equations of motion can be expressed in first-order form
$(i\partial_{z}+\omega+\mathcal{M})\psi=0.$ (S16)
In the mixing matrix, we have neglected the absorption effects of photons, the contribution of the cosmic microwave background’s energy density to the dispersion relation, and the QED vacuum birefringence effect.
Assuming a small DM virial velocity $v_{\mathrm{DM}}\ll 1$, we model the background field as
$\bar{a}(t,z)=m_{a}^{-1}\sqrt{2\rho_{\text{DM}}}\cos(m_{a}t+\varphi_{a}),$ (S17)
where $\rho_{\text{DM}}$ is the DM density and $\varphi_{a}$ is the phase of the axion background. This leads to an oscillating birefringence term:
$\Delta_{F}=-i\Delta_{\text{DM}}\sin(m_{a}t+\varphi_{a}^{\prime}),$ (S18)
where the amplitude is $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{(1+\mathcal{F}^{2})\rho_{\text{DM}}/2}$ and the phase is $\varphi_{a}^{\prime}=\varphi_{a}+\arctan\mathcal{F}$. The factor $\mathcal{F}(z)=m_{a}^{-1}d\ln\rho_{\text{DM}}/dz$ can be used to describe possible corrections from spatial DM gradients, but will be neglected in the following. For fuzzy DM, $\mathcal{F}(z)$ is less than the virial velocity of DM. Along the photon trajectory we take the eikonal propagation approximation $t\simeq z$ (with $c=1$).
II Effective Two-Level systems
Applying $\psi(z)=e^{i\omega z}\Psi(z)$, the reduced Schrödinger-like equation can be written as
$i\frac{d}{dz}\Psi(z)=\mathcal{H}(z)\Psi(z),\quad\mathcal{H}(z)=-\mathcal{M}(z),$ (S19)
where $\mathcal{M}$ is the mixing matrix including the oscillating birefringence term. To solve the dynamics analytically, we perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ to eliminate the rapid oscillations in the photon sub-block. The effective Hamiltonian becomes
$\mathcal{H}^{\prime}(z)=U^{\dagger}\mathcal{H}U-iU^{\dagger}\frac{dU}{dz}.$ (S20)
Define $\phi(z)\equiv m_{a}z+\varphi_{a}$, the photon sub-sector of $\mathcal{H}$ can be decomposed using Pauli matrices:
$\mathcal{H}_{\gamma}(z)=-\Delta_{\gamma}\,\mathbb{I}_{2}-\delta\Delta_{\gamma}\,\sigma_{z}-\Delta_{\text{DM}}\sin\phi\,\sigma_{y},$ (S21)
where $\Delta_{\gamma}=(\Delta_{\perp}+\Delta_{\parallel})/2$ and $\delta\Delta_{\gamma}=(\Delta_{\perp}-\Delta_{\parallel})/2$. We introduce a time-dependent unitary rotation acting in the photon subspace
$u(z)=\exp\!\big[-i\theta(z)\sigma_{y}\big],\quad\theta(z)=\alpha\cos\phi(z).$ (S22)
The transformation matrix is
$U(z)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}$. (S23)
The term $iu^{\dagger}\partial_{z}u$ can be calculated as
$iu^{\dagger}\partial_{z}u=-\alpha m_{a}\sin\phi\;u^{\dagger}\sigma_{y}u=-\alpha m_{a}\sin\phi\;\sigma_{y},$ (S24)
where the final step utilises the fact that $\sigma_{y}$ commutes with itself. To cancel the $\Delta_{\text{DM}}\sin\phi\,\sigma_{y}$ term in the Hamiltonian, we set $\alpha=\Delta_{\text{DM}}/m_{a}$, yielding the effective Hamiltonian
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}+\delta\Delta_{\gamma}\cos(2\theta)&\delta\Delta_{\gamma}\sin(2\theta)&\Delta_{B}\sin\theta\\[6.0pt] \delta\Delta_{\gamma}\sin(2\theta)&\Delta_{\gamma}-\delta\Delta_{\gamma}\cos(2\theta)&\Delta_{B}\cos\theta\\[6.0pt] \Delta_{B}\sin\theta&\Delta_{B}\cos\theta&\Delta_{a}\end{pmatrix}$. (S25)
When $|\delta\Delta_{\gamma}|\ll m_{a}$ and $\Delta_{\mathrm{DM}}$, the phase evolution rate due to QED birefringence becomes negligible compared to the DM modulation frequency. In this regime, the photon states are effectively degenerate for axions, allowing the QED contribution to be safely neglected. The transformed Hamiltonian simplifies to
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$. (S26)
For $\Delta_{\mathrm{B}}\ll|\Delta_{\gamma}-\Delta_{a}|$, the dynamics can be analyzed by treating each photon mode-axion pair independently under the rotating-wave approximation. The two-state systems are governed by couplings $G^{(1)}=\Delta_{\mathrm{B}}\sin\theta$ and $G^{(2)}=\Delta_{\mathrm{B}}\cos\theta$. Expressing $\sin\theta$ and $\cos\theta$ via the Jacobi–Anger identity
$e^{i\alpha\cos\phi}=\sum_{\ell=-\infty}^{\infty}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi},$ (S27)
we expand $G^{(1,2)}$ into Fourier harmonics:
$G^{(1)}(z)$ $=\frac{\Delta_{\mathrm{B}}}{2i}\sum_{\ell}[1-(-1)^{\ell}]i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi},$ (S28)
$G^{(2)}(z)$ $=\frac{\Delta_{\mathrm{B}}}{2}\sum_{\ell}[1+(-1)^{\ell}]i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}.$ (S29)
Thus, each photon mode couples to the axion through a series of sidebands with frequencies $\ell m_{a}$. $G^{(1)}$ and $G^{(2)}$ carry odd and even harmonics, respectively. We define the relevant coupling as
$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z},$ (S30)
with
$g_{\ell}^{(1)}$ $=\frac{[1-(-1)^{\ell}]i^{\ell-1}e^{i\ell\varphi_{a}}}{2}J_{\ell}(\alpha)\Delta_{\mathrm{B}},$ (S31)
$g_{\ell}^{(2)}$ $=\frac{[1+(-1)^{\ell}]i^{\ell}e^{i\ell\varphi_{a}}}{2}J_{\ell}(\alpha)\Delta_{\mathrm{B}}.$ (S32)
Focusing on a single photon mode, the coupled equations become:
$-i\frac{d}{dz}\Phi_{\gamma}=\Delta_{\gamma}\Phi_{\gamma}+G(z)\Phi_{a},$ (S33)
$-i\frac{d}{dz}\Phi_{a}=G(z)\Phi_{\gamma}+\Delta_{a}\Phi_{a}.$ (S34)
Moving to the interaction picture via $\Phi_{\gamma}=\tilde{\Phi}_{\gamma}e^{i\Delta_{\gamma}z}$, $\Phi_{a}=\tilde{\Phi}_{a}e^{i\Delta_{a}z}$, we obtain
$-i\frac{d}{dz}\tilde{\Phi}_{\gamma}=\sum_{\ell}g_{\ell}e^{-i\Delta_{\ell}z}\tilde{\Phi}_{a},$ (S35)
$-i\frac{d}{dz}\tilde{\Phi}_{a}=\sum_{\ell}g_{\ell}^{*}e^{i\Delta_{\ell}z}\tilde{\Phi}_{\gamma},$ (S36)
with
$\Delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}.$ (S37)
Resonance occurs when the phase oscillation is slow, i.e., when the momentum mismatch is compensated by the axion mass harmonics. Retaining the resonant harmonic $\ell=n$ that satisfies $\Delta_{n}\approx 0$, we obtain the Rabi oscillation dynamics with an effective Rabi frequency
$\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\Delta_{n}/2)^{2}}.$ (S38)
The conversion probability is given by:
$P_{\gamma\to a}=\frac{|g_{n}|^{2}}{\Omega_{\rm R}^{2}}\sin^{2}\big(\Omega_{\rm R}\,z\big).$ (S39)
Additionally, we discuss the choice of transformations for diagonalizing the Hamiltonian $\mathcal{H}$. In general, Floquet theory can be directly applied to the three-level system. However, to isolate the dominant resonant dynamics and obtain a transparent physical picture, we diagonalize the photon subspace and reduce the problem to effective two-level photon–axion systems. An intuitive choice is to transform to the circular polarization basis using
$U_{C}(z)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i&0\\ 1&-i&0\\ 0&0&1\end{pmatrix}$. (S40)
The Hamiltonian is given by
$\mathcal{H}_{C}^{\prime}=-\begin{pmatrix}\Delta_{+}&\delta\Delta_{\gamma}&\frac{\Delta_{\mathrm{B}}}{\sqrt{2}}\\ \delta\Delta_{\gamma}&\Delta_{-}&\frac{i\Delta_{\mathrm{B}}}{\sqrt{2}}\\ \frac{\Delta_{\mathrm{B}}}{\sqrt{2}}&-\frac{i\Delta_{\mathrm{B}}}{\sqrt{2}}&\Delta_{a}\end{pmatrix}$ (S41)
where $\Delta_{\pm}=\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi$. Neglecting the QED birefringence effect, the evolution for the photon state can be formally written as
$\Phi_{\pm}$ $\simeq\exp\left[-i\int^{z}dz^{\prime}(\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi(z^{\prime}))\right]\tilde{\Phi}_{\pm}$ (S42)
$=e^{-i\Delta_{\gamma}z}e^{\pm i\alpha\cos\phi}\tilde{\Phi}_{\pm}$. (S43)
Repeating the Fourier harmonic expansion as above, one would also obtain the same resonance condition and Rabi oscillations. Although axion-induced frequency modulation has been noted in earlier studies [McDonald:2019wou], it is often incorporated as a slowly varying phase or an effective modification of the refractive index. Such treatments effectively average over the harmonic structure of the modulation and therefore do not capture the Floquet sidebands responsible for resonant mode mixing.
The transformations $U(z)$ and $U_{C}$ offer two distinct yet complementary physical interpretations. The transformation $U(z)$ represents a co-rotating frame perspective. By shifting into a basis that rotates alongside the axion-induced birefringence effect, we effectively eliminate the rapid oscillations of the off-diagonal mixing terms. In this frame, the interaction is understood through the lens of geometric phase, where the coupling to the axion arises from the mismatch between the photon’s polarization drift and the background field’s oscillation frequency.
The circular basis $U_{C}$ provides a frequency modulation perspective. Here, the axion field acts as a dynamic medium that oscillates the refractive indices of the left- and right-handed circular polarizations in opposite phases. The resonance condition is explained by the Jacobi-Anger expansion, where the carrier frequency of the photon acquires sidebands at integer multiples of the axion mass $m_{a}$. While both methods yield identical Rabi frequencies for the resonant transition, the circular basis is often more robust when incorporating higher-order corrections, such as QED birefringence, as it treats the two helicity states as the fundamental vacuum eigenmodes.
III Coherent Mode Mixing vs. Parametric Instability
In the literature on axion electrodynamics, it is well established that an oscillating axion background can induce a parametric instability in the electromagnetic field. This is often formulated using the Mathieu equation and interpreted as the stimulated decay of axions into photons. Superficially, this resembles the propagation equation, Eq. (S12), utilized in this work, where the mixing matrix $\mathcal{M}(z)$ also contains a periodic contribution from the axion DM background. In this appendix, we clarify that these two phenomena, despite originating from the same Lagrangian, describe physically distinct regimes and yield qualitatively different results.
III.1 Common Origin
Both formalisms derive from the standard axion–photon Lagrangian Eq. (S1). Consider a classical, coherently oscillating axion background
$a(t)\sim a_{0}\cos(m_{a}t).$ (S44)
Maxwell’s equations acquire an explicit time dependence, characterized by Eqs. (S6) to (S8). However, the subsequent reduction of these equations depends critically on the physical boundary conditions and the kinematic regime of interest.
III.2 Regime I: Parametric Instability (The Mathieu Limit)
The analysis of axion-induced parametric instability typically focuses on the spontaneous production and amplification of photon modes from the vacuum (or a low-occupation state). In this regime, one must retain the full second-order time dynamics, including both positive- and negative-frequency components of the field momentum modes:
$\mathbf{A}(\mathbf{x},t)\sim A_{\pm}(t)\,e^{i\mathbf{k}\cdot\mathbf{x}},\qquad A_{\pm}(t)=c(t)e^{-i\omega t}+c^{*}(t)e^{+i\omega t}.$ (S45)
For circular polarizations, the equation of motion reduces to a Mathieu equation:
$\ddot{A}_{\pm}+\omega^{2}\left[1\pm\epsilon\cos(m_{a}t)\right]A_{\pm}=0,$ (S46)
where $\epsilon\sim g_{a\gamma}a_{0}m_{a}/\omega$. The defining feature of this regime is that the periodic axion background modulates the effective frequency squared of the oscillator. This modulation couples the positive-frequency ($e^{-i\omega t}$) and negative-frequency ($e^{+i\omega t}$) modes. When $m_{a}\simeq 2\omega$, the system enters an instability band where the field amplitude grows exponentially. This corresponds to the stimulated decay of the axion field into photon pairs, a genuine parametric instability where the background acts as an energy pump.
III.3 Regime II: Coherent Mode Mixing (The Propagation Limit)
In contrast, this work investigates the propagation of a pre-existing electromagnetic wave with a fixed carrier frequency $\omega$ and well-defined direction. Employing the slowly varying envelope approximation, we express the field as a slowly evolving amplitude modulating a rapid oscillatory phase. Crucially, this procedure effectively integrates out backward-propagating and negative-frequency modes. The resulting dynamics are governed by Eq. (S12). While the axion background still introduces a periodic modulation in the mixing matrix $\mathcal{M}(z)$, its physical role changes fundamentally:
  • The modulation acts as a linear, Hermitian coupling between forward-propagating modes (e.g., photon polarization states).
  • It does not couple positive- and negative-frequency solutions.
Consequently, the system does not exhibit Mathieu-type exponential instabilities. Instead, it displays Floquet resonances—enhanced but bounded mode conversion—when the eigenvalue difference matches the modulation frequency: $\lambda_{i}-\lambda_{j}=nm_{a}$.
III.4 Theoretical Synthesis
The distinction between these two phenomena can be understood by analogy to quantum optical systems.
  • Parametric Instability (Regime I) is analogous to a driven harmonic oscillator. A periodic modulation of the oscillator’s parameters leads to exponential growth of the mode amplitude within instability bands. In quantum terms, this represents parametric amplification, or particle creation from the vacuum.
  • Floquet Mixing (Regime II) is analogous to a driven multi-level system, such as Rabi oscillations in atomic physics. The relevant degrees of freedom are discrete internal states, and the periodic background induces transitions among them. Probabilities oscillate but remain bounded (unitary evolution).
From a formal Floquet theory perspective, both systems involve linear differential equations with periodic coefficients. However, the spectral implications differ:
  • In the Mathieu case (Instability), the Floquet quasi-energies can acquire imaginary parts, signaling an unbounded instability.
  • In the Propagation case (Mixing), the quasi-energies remain real. Resonances here correspond to avoided crossings in the quasi-energy spectrum, maximizing the mixing angle between states rather than the total energy of the system.
Thus, while both phenomena arise within the framework of periodically driven systems, the propagation formalism employed in this work isolates the physics of coherent state mixing, explicitly distinct from the particle-production physics of parametric instability.
IV Numerical analysis
To numerically verify the analytical framework, we solve the Eq. (S16) incorporating the mixing matrix $\mathcal{M}$ for a concrete astrophysical environment. We model the relativistic jet of blazar 3C 279, focusing on the unpolarized optical R-band photons ($E_{\gamma}\approx 1.9\,\mathrm{eV}$) with a representative parameter set: an axion mass $m_{a}=10^{-22}\,\mathrm{eV}$ and a coupling constant $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$. The radiation zone, located $r_{0}=0.5\,\mathrm{pc}$ from the central engine, is characterized by a magnetic field $B_{0}=0.5\,\mathrm{G}$ and electron density $n_{e,0}=5\times 10^{3}\,\mathrm{cm}^{-3}$ [A:2025wjt]. The jet is modeled as a conical outflow with a Doppler factor $\delta\simeq 13$ [A:2025wjt], assuming scaling profiles $B(r)\propto r^{-1}$ and $n_{e}(r)\propto r^{-2}$ over a total length $L_{\rm jet}=1\,\mathrm{kpc}$. To capture the fine-grained evolution of the mixing process, the calculation is performed over a domain size ranging from $5\times 10^{-7}\,\mathrm{pc}$ to $10^{-3}\,\mathrm{pc}$. To isolate and illustrate the resonance structures dictated by the plasma profile, we adopt a uniform DM density $\rho_{\rm DM}=1\,\mathrm{GeV\cdot cm^{-3}}$ ($2.63\times 10^{7}\,M_{\odot}\mathrm{\cdot kpc^{-3}}$).
The numerical results of the photon–axion conversion probability in Fig. 1 demonstrate resonant conversion exactly at the locations predicted by Eq. (10), while the behavior reduces to the adiabatic limit away from resonance. This excellent agreement confirms that the Floquet analysis accurately describes the physics of the driven system. The conversion probability exhibits a distinct step-like growth, indicating that resonant conversion occurs within narrow spatial windows where the phase-matching condition is satisfied. The resonance positions depend solely on the jet’s plasma profile and the axion mass, independent of the DM density $\rho_{\rm DM}$. Thus, while a realistic non-uniform DM halo would modulate the amplitude of each ”step” according to local $\rho_{\rm DM}$ and $B$ values, the qualitative structure of multiple resonances remains unchanged.
The spatial width of the parametric resonance $L_{\text{res}}$ can be estimated as $\left|4g_{n}/(\nabla_{r}\Delta_{n})\right|_{r_{\text{res}}}$. For the jet model considered here, the resonance width is approximately $2(nm_{a})^{-1}g_{a\gamma}B_{0}J_{n}(\alpha)r_{0}$. For the parameters used in Fig. 1, we obtain $L_{\text{res}}\sim 3\,\text{pc}$, in agreement with the numerical results. This finite width confirms that the resonance is robust against density fluctuations on scales smaller than $L_{\text{res}}$. Given that $L_{\text{res}}$ is two orders of magnitude smaller than the total jet length ($\sim\mathrm{kpc}$), plasma density variations within each resonance region are negligible. This separation of scales ensures that the parametric resonance remains robust even in highly inhomogeneous jet environments, where the global density gradient merely serves to sequentially satisfy discrete resonance conditions.
The parametric resonance relies on the coherent oscillation of the axion field. The validity of this assumption depends on the coherence time of the DM field, $\tau_{\mathrm{coh}}\sim(m_{a}v^{2})^{-1}$, compared to the photon’s transit time through the resonance region, $t_{\mathrm{cross}}\sim L_{\mathrm{res}}$. For ultralight axions considered here with virial velocity $v\sim 10^{-3}$, the coherence time is $\tau_{\mathrm{coh}}\sim\mathcal{O}(10^{5})\,\text{years}$. In contrast, the spatial width $L_{\mathrm{res}}\sim 3\,\mathrm{pc}$ corresponds to a transit time $t_{\mathrm{cross}}\approx 10\,\text{years}$. The condition $t_{\mathrm{cross}}\ll\tau_{\mathrm{coh}}$ ensures that the DM background acts as a highly coherent driving force during the resonant interaction. Moreover, the fractional frequency spread $\delta m_{a}/m_{a}\sim 10^{-6}$ induces a negligible broadening of the resonance layer compared to its physical width determined by the jet geometry and coupling strength.
Figure S1: ECDF of the photon-to-axion conversion probability $P_{\gamma\to a}$. The color coding indicates various turbulence correlation lengths $L_{fluc}$ for electron density (from $10^{-3}$ to $1\text{ pc}$), while solid and dashed lines distinguish between models with and without DM background, respectively. Each curve is derived from a sample size of $N=1000$. Note that the x-axis is presented in log-scale.
To assess the robustness of the resonance mechanism, we introduce 50% stochastic electron density fluctuations across turbulence scales ranging from $L_{\rm fluc}=10^{-3}$ to $1$ pc. As shown in Fig. S1, the mean conversion probability in the DM background remains consistently 1-2 orders of magnitude higher than the standard baseline without DM, regardless of the correlation length. While small-scale fluctuations ($L_{\rm fluc}\ll L_{\rm res}$) are effectively self-averaged, larger-scale variations ($L_{\rm fluc}\sim 1$ pc) increase statistical variance by inducing local dephasing. Nevertheless, the empirical cumulative distribution function reveals that even the lower quantiles of these distributions significantly surpass the standard adiabatic case.
This resilience stems from the ”hard-driving” nature of the Floquet resonance, where the global coherence of the axion oscillation frequency $m_{a}$ exerts a persistent periodic force that facilitates mode-mixing despite environmental noise. Even when the resonance condition is fragmented by medium-scale turbulence, the collective accumulation of these imperfect resonances ensures a statistically robust conversion probability.
Standard astrophysical emission mechanisms in blazars typically yield negligible circular polarization. However, parametric resonance breaks the degeneracy between left- and right-handed circular polarization states through the parity-violating coupling. This induces a non-zero net circular polarization, whose sign and magnitude are stochastic and depend on the phase $\varphi_{a}$ of the axion background at the resonance onset.
We apply the formalism to constrain the axion–photon coupling $g_{a\gamma}$ using optical polarimetry of 3C 279. The DM density is modeled using a fuzzy DM profile, which consists of a solitonic core and an outer NFW envelope [Navarro:1995iw, Schive:2014dra] with the total mass $M_{h}\sim 10^{13}M_{\odot}$ [Ferrarese:2002ct]. The core density profile is adopted as $\rho_{c}[1+0.091(r/r_{c})^{2}]^{-8}$ with the core radius $r_{c}\simeq 1.6\,\mathrm{kpc}\left(m/10^{-22}\,\mathrm{eV}\right)^{-1}\left(M_{\rm h}/10^{9}\,M_{\odot}\right)^{-1/3}$ and density $\rho_{c}\simeq 2.9\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-3}(m/10^{-22}\,\mathrm{eV})^{2}(M_{\rm h}/10^{9}\,M_{\odot})^{4/3}$ [Schive:2014hza]. The transition radius between the two regions is taken to be $r_{t}\simeq 3r_{c}$. The concentration parameter of the NFW halo is determined using the concentration–mass relation at redshift $z=0.5$ [Dutton:2014xda]. This configuration yields negligible spatial gradients with $\mathcal{F}\lesssim 10^{-3}$ .
Figure S2: KDE-estimated probability density of induced circular polarization for various $m_{a}$ ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$), obtained by uniformly sampling the initial axion phase. Shaded areas denote density profiles computed using Gaussian kernels with Scott’s rule bandwidths.
For each parameter pair $(m_{a},g_{a\gamma})$, we solved the equations of motion 1,000 times, using equally spaced values between 0 and $2\pi$ for the initial axion phase $\varphi_{a}$ and assuming a $30\%$ initial linear polarization [Liodakis:2021els], then we obtained a well-resolved ensemble of induced circular polarization. In Fig. S2, the resulting probability density of the degree of circular polarization $\Pi_{C}$ is visualized using the kernel density estimation. By comparing the variance of this predicted signal with characteristic observational uncertainties, we set conservative consistency limits on the axion–photon coupling $g_{a\gamma}$ for the mass range $10^{-23}-10^{-20}\,\mathrm{eV}$ in Fig. 2.